b.Using the fact that PX=12nWKA94cg+6jgAAAABJRU5ErkJggg==, compute the expected value E(X) and standard deviation (σ) of the number of pages for the population of 2n WaUYCizxc8CSilYCdK6gHfDBYB2uTgGWGhOAyqy4books.

QUESTION 1 (39 MARKS) Suppose you choose 2 books that you are planning to read (the books selected must contain both Malay and English). Thus, there will be a total of 2n number of books, where n is the number of team members. Assume that this 2n is the size of population that is approximately normally distributed with mean and standard deviation ơ a) List the number of pages for each of the 2n books as in Table 1. Let X be the number of pages for the selected books. Send the completed table in Excel file to your lecturer Table 1: Team Data Book ID Book Title Frozen 32 Budak mekanikal 40 73 You to me Demi masa 80 43 Positive making guide Hürah 60 (4 marks) b) Using the factx compute the expected value EX) and standard deviation (a) of the number of pages for the population of 2n books (7 marks)
Answer #1
  • Expected value, E(X) = \sum X.P(X)

    P(X) is the same for all the values of X i.e. P(X) = 1/2n = 1/6

    So, expected value becomes: E(X) = P(X)ΣΧ=6(32 + 40 + 73 + 80+43 + 60)

    ==> E(X) 328/6-54.667

    --------------------------------------------------------------------------------------------------------------------------

    Standard deviation, VPX).(X- E(X))

    1 (32 54.667)2(40-54.667)2 (73 54.667)2 (80 54.667)(43 54.667)2(60- 54.667)2)

    ==> σ 17.660